I need help remembering what the standard formula is. I believe the slope intercept formula is.. y = mx+b and point slope formula is .. y-y1= m(x-x1). Are these two correct? The specific question I having trouble with is: Given the following, write an equation in standard form. The line has y-intercept of 5 and a slope of 2. If you could be of any help I would most appreciate it!
Thank you!

Answer by Mike(39)

By standard form they mean: ax + by = c. Point slope form is: y-a = m(x-b) and slope-intercept form you said correctly, as it is y=mx+b.
In answering this question you must remember the following - when you are asked to write something in standard form, 'a' must be positive and all values in the equation must be whole numbers.
In starting to answer this question, I would first write it as follows:
y = 2x + 5, and then:
2x - y = -5
It's as easy as that.
Hope that helped.
Regards,
-Mike